What is the ? In the following equation
20+5×(?+6)-12÷2=64

Answer by Hawksfan(61)

20 +5x(?+6)-12/2 = 64
Because * and / have preference over + and -, we need to divide 12/2.
first to get
20 + 5*(? +6) -6 =64
combine 20 -6 to get 5(?+6) +14 = 64
subtract 14 5(?+6) =50
Divide by 5 ? + 6 = 10

subtract 6 ? = 4

check 64= 64